# LASER: Lip Landmark Assisted Speaker Detection for Robustness

### v1

@@ 4.3.1 Integrating Lip Track Encoding to Visual Features
As discussed in Section 3.2, we integrate the aggregated lip track encoding (LTE) into the visual features produced by the 3D-CNN. We compare this approach with integrating LTE directly with the RGB face frames before the 3D-CNN. As shown in Table 2, integrating LTE before the 3D-CNN leads to a significant performance drop. We speculate that this is due to the limited number of channels in the RGB image, where integrating LTE with additional channels may interfere with the feature extraction process of the original face frame. Thus, we opt to integrate LTE into the visual features output by the 3D-CNN.
@@
Table 4: Ablation study on lip track integration with different stages of ResNet. Setting $S=4$ and integrating lip track encoding before the first stage of ResNet results in the best performance.
We further investigate integrating LTE to different stages within ResNet. ResNet models have 4 internal stages with different downsampling strides and we investigate the choice of integration before each of the four stages. As shown in Table 4, integrating LTE to features before entering the first stage (i.e, the output feature from 3D-CNN) achieves the overall best performance when evaluated with noise audio tracks and integrating it to subsequent stages within ResNet consistently decreases the performance. Therefore, we choose to integrate LTE to the visual features output by 3D-CNN (i.e, before the first stage of ResNet).
4.3.2 Number of Lip Track Encoding Channels
Recall that we aggregate the initial sparse lip track encodings from $K=82$ to $S=4$ dense feature maps through our aggregation process to produce LTE. Here, we examine the sensitivity of the choice of $S$. As shown in Table 4, the choice of $S$ involves a tradeoff: setting $S$ too low overly condenses the information, while setting $S$ too high results in feature maps that remain sparse. Thus, we select $S=4$, which provides an optimal balance.
@@ 4.3.2 Number of Lip Track Encoding Channels
Table 5: Ablation study on consistency loss weight. Setting $\lambda_{c}=1$ achieves the best results while decreasing $\lambda_{c}$ consistently degrades the performance.
Figure 5: Qualitative results of LASER, TalkNCE, LoCoNet on two unsynchronized videos. We create non-speaking scenarios by swapping the audio tracks of two videos where the same person is speaking with similar camera angles. The red box means the model predicts not-speaking. The green box means the model predicts speaking. The yellow circle means the model makes a wrong prediction.
4.3.3 Contribution of Consistency Loss
The consistency loss is a key component of LASER designed to address the issue of missing lip track data during test time. Our analysis reveals that the face landmark detector struggles to predict lip landmarks for approximately 15% of the testing face tracks in the AVA-ActiveSpeaker dataset. As shown in Table 3, models trained with the consistency loss achieve a significant improvement when evaluated with both regular and noisy audio inputs. Moreover, when using consistency loss, our models deliver comparable results whether or not lip track encoding is available at test time. This indicates that the visual encoder can successfully capture lip movements through consistency loss, even without the lip track encoding during inference. Consequently, at test time, our model does not rely on the face landmark detector for lip track prediction, which helps reduce the inference latency.
4.3.4 Weight of Consistency Loss
We also study the sensitivity of the weight of consistency loss $\lambda_{c}$ where we range the value from 0.2 to 1.0. As shown in Table 5, $\lambda_{c}=1$ achieves the best performance whereas decreasing the loss weight consistently hurts the model performance. Thus, we set $\lambda_{c}=1$ in all of our experiments.
4.4 Qualitative Results
We provide qualitative comparisons of active speaker detection with real-world videos. Specifically, we collect two video clips with the same person speaking from the same camera angles yet different content in the speech. We then reverse the audio track of the two video clips to create unsynchronization between face tracks and audio tracks.
As shown in Figure 5, the baseline LoCoNet method experiences challenges when predicting non-speaking in such a challenging scenario whereas our LASER succeeds to make correct predictions in this scenario. The results demonstrate that our LASER is more robust on challenging real-world scenarios.
5 Conclusion
In this work, we introduced LASER for active speaker detection. LASER stands apart from existing methods by explicitly guiding the model to focus on lip landmarks, rather than other raw pixel based facial expressions, enhancing the model’s ability to accurately detect active speakers and improving its sensitivity to audiovisual synchronization. Both quantitative and qualitative results demonstrate that LASER delivers stronger performance compared to the state-of-the-art, especially when visual frames and audio tracks are not synchronized, and exhibits more robust results in real-world video scenarios.

### v2

@@ A.4.1 ResNet vs 3D CNN
As discussed in Section 3.2, we integrate the aggregated lip track encoding (LTE) into the visual features produced by the 3D-CNN. We compare this approach with integrating LTE directly with the RGB face frames before the 3D-CNN. As shown in Table B, integrating LTE before the 3D-CNN leads to a significant performance drop. We speculate that this is due to the limited number of channels in the RGB image, where integrating LTE with additional channels may interfere with the feature extraction process of the original face frame. Thus, we opt to integrate LTE into the visual features output by the 3D-CNN.
A.4.2 Integrating Lip Track Encoding to Visual Features
@@
Table A: Ablation study on lip track integration with different stages of ResNet. Setting $S=4$ and integrating lip track encoding before the first stage of ResNet results in the best performance.
Layer Regular Audio Noise Audio
3D CNN 92.6 86.6
ResNet 95.1 86.7
Table B: Ablation study on lip track encoding integration. Integrating lip track encoding before 3D-CNN layer (i.e, on RGB frames) results in suboptimal performance.
We further investigate integrating lip landmark encoding (LTE) at different stages within ResNet, which has four internal stages with varying downsampling strides. Our ablation study (Table A) shows that integrating LTE before the first stage—using features from the 3D-CNN output—yields the best performance under noisy audio conditions, while later-stage integration degrades performance. Additionally, we examine the sensitivity of the aggregation factor $S$ when condensing sparse lip track encodings from $K=82$ to $S$ dense feature maps. We find that $S=4$ achieves the best balance, avoiding both excessive compression and sparsity.
@@
Table C: Ablation study on consistency loss weight. Setting $\lambda_{c}=1$ achieves the best results while decreasing $\lambda_{c}$ consistently degrades the performance.
A.4.3 Weight of Consistency Loss
We also study the sensitivity of the weight of consistency loss $\lambda_{c}$ where we range the value from 0.2 to 1.0. As shown in Table C, $\lambda_{c}=1$ achieves the best performance whereas decreasing the loss weight consistently hurts the model performance. Thus, we set $\lambda_{c}=1$ in all of our experiments.